# Navigating the Reality Gap: Privacy-Preserving Adaptation of ASR for Challenging Low-Resource Domains

### v2

@@ 7 Limitations and Future Directions
While our framework demonstrates significant potential for deploying ASR in resource-constrained environments, several limitations remain:
@@
    Acoustic-Only Training: Our primary training framework focuses exclusively on adapting the acoustic model (AM). While our validation "spot check" confirmed that even a domain-specific LM is insufficient without acoustic adaptation (Section 5), we did not integrate the LM into the continuous training loop (e.g., via shallow fusion during replay). Capturing the synergy between acoustic and linguistic adaptation during training remains a key area for future work.
@@
    Reliance on Supervision: Our "continual learning" assumption relies on the availability of a stream of corrected transcripts (e.g., from medical professionals correcting dictations). In scenarios where such feedback is sparse, delayed, or noisy, the adaptation rate would likely degrade.
@@
    Static vs. Dynamic Adaptation Scheduling: Our analysis suggests that an optimal adaptation strategy would ideally employ Experience Replay in the initial phases to rapidly converge to the target domain, followed by a transition to aggressive regularization as the model approaches the “Adaptation Wall” to prevent stability collapse. However, because this wall is inherently dependent on the language, model, and domain, its onset is difficult to predict. Further research into the fundamental limits of adaptation in continual learning, particularly through the integration of language models that provide linguistic context will be critical for refining these strategies.
@@
    Language Family Scope: Our experiments were conducted on Hindi and its rural dialects. While we hypothesize the findings apply to other Indo-Aryan languages, the efficacy of this specific replay strategy on tonal languages or those with fundamentally different acoustic structures (e.g., Dravidian languages) remains to be validated.

### v3

@@ 6 Conclusion
7 Privacy, Ethics, and Clinical Safety Considerations
The deployment of ASR in frontline healthcare settings introduces unique ethical and safety considerations which our framework explicitly addresses.
Data Governance and Residency:
Our primary focus on on-device adaptation is a direct response to healthcare data residency requirements. By ensuring that raw patient audio never leaves the local environment, we mitigate the risk of centralized data breaches and maintain compliance with emerging data protection laws in various jurisdictions (nawaz2019privacy; rieke2020future).
Clinical Safety and Error Risks:
While our framework significantly reduces WER, automated transcription in medical contexts is never risk-free. ASR errors in medically salient keywords (e.g., negations or dosage numbers) can lead to serious diagnostic errors. We advocate for a Clinician-in-the-Loop model, where our adaptive system serves as a draft-generation tool that reduces manual workload while maintaining human oversight for final clinical validation.
Algorithmic Bias and Equity:
By prioritizing adaptation on rural dialects and low-bandwidth telephony, this work contributes to healthcare equity. Standard ASR models often fail on marginalized populations due to accent mismatch; our framework allows localized clinics to "correct" these biases in real-time by adapting specifically to their local community’s speech patterns.
@@ 8 Limitations and Future Directions
While our framework demonstrates significant potential for clinical deployment, several limitations remain:
@@
    Metrics as a Clinical Proxy: Our evaluation relies on WER and CER, which are standard in ASR but do not explicitly weight the clinical significance of errors. An error in a patient’s name may be less critical than an error in a medication dosage; developing medical-concept-aware metrics remains an essential future direction.
@@
    Reliance on Clinician Supervision: Our continual learning assumption assumes a stream of corrected transcripts (e.g., from medical professionals). In under-staffed rural clinics, this supervision may be sparse or delayed, necessitating future work into uncertainty-aware or semi-supervised adaptation.
@@
    Acoustic-Only Training: While we validated the necessity of acoustic adaptation, we did not integrate Language Models (LMs) into the training loop (e.g., via shallow fusion during replay). Capturing the synergy between acoustic and linguistic adaptation is critical for handling specialized medical terminology.
@@
    Language and Dialect Scope: While we focus on rural Hindi dialects, the efficacy of our replay strategy on tonal or Dravidian languages remains to be validated.